OpenStudy (anonymous):

A particle moving in a straight line has an acceleration given by a(t)=2t. The initial velocity of the particle is 2 cm/sec. How far does the particle move between t =1and t = 2 seconds?

OpenStudy (anonymous):

\[X = \frac{ 1 }{ 2 }at^2+Vi*t +X i \] so X-X i = \frac{ 1 }{ 2 }at^2+Vi*t